OpenStudy (mirandaregina):

I need help finding the y-values of the function y=sin(x+pi/2). Here's what I have so far: Amp = 1, Horizontal Shift = pi/2 to the left, Vertical Shirt = none, Period = 2pi/2 = pi. Then I divided the period (pi) by 4 to get pi/4, and translated the x-values of the parent graph y=sinx to pi/4, pi/2, 3pi/4, pi, and 5pi/4. How do I find the y-values from here?

OpenStudy (shubhamsrg):

Why would there be any change in the period ?

OpenStudy (mirandaregina):

@shubhamsrg no, the period remains that same, but I divided 4 by the period. So 0 + pi/4 = pi/4, pi/4 + pi/4 = 3pi/4 and so on. I graphed y=sin2(x+pi/2) on the website desmos, and I know that my x-values are correct, i just don't know how to get the y-values, which are apparently -1, 0, 1 0, -1 for the first five critical points.

OpenStudy (mirandaregina):

@yayanieves

OpenStudy (anonymous):

Sorry cant help you. Tryna figure my own question out. Sorry.

OpenStudy (mirandaregina):

@yayanieves I understand, thanks anyway!